That is 23.0% where the cursor is.
Correct Significant Figures Feedback: Your answer 949.21 m was either rounded differently or used a different number of significant figures than required for this part If you need this result for any later calculation in this item, keep all the digits and round as the final step before submiting your answer. Constants The food calorie, equal to 4186 J, is a measure of how much energy is released when food is metabolized by the body. A certain brand of fruit and-cereal bar contains 160 food calories per bar Part B it,as is typical only,0% (Note: In this and al other problems, we are assuming that 100 % of the food calories that are eaten are absorbed and used by the body. This is actually not true. A persons metabolic efficiency is the percentage of calories eaten that are actualy used, the rest are eliminated by the body. Motabolic efficiency varios considerably from person to person,) of the food calories go into mechanical energy, what would be the answer to part (a)? Submit Request Answer Provide Feedback Next > Previous Next w
Answer #1

Given that,

23.0% of the food calories go into mechanical energy

we are assuming that 100 % of the food calories that are eaten are absorbed and used by the body

Energy = 0.23 * 564,300 = 129,789

Work = Energy
70 * 9.8* h = 129,789
h = 189.2 meters

Please rate it..............